# On $p$-adic Asai $L$-functions of Bianchi modular forms at non-ordinary primes and their decomposition into bounded $p$-adic $L$-functions

### v2

@@ 4.2.1. Another description of weight $2$ Asai-Eisenstein elements
[21, Lemma 4.5] The pushforward of $\prescript{}{c}{C}_{m^{2}N}$ along the map $Y_{\mathbb{Q},1}(m^{2}N)\to Y_{\mathbb{Q}}(m,mN),$ given by $z\mapsto mz$ on $\mathbb{H}$, is $\prescript{}{c}{C}_{mN}$.
@@
[21, Proposition 4.4] Let $a\in\mathcal{O}_{F}$ be a generator of $\mathcal{O}_{F}/(p\mathcal{O}_{F}+\mathbb{Z})$. We have
@@ 4.3.2. Relation to the weight $2k$ Eisenstein series
From [21, Proposition 5.2], for an integer $k\geq 0$, we know that there exists a class $\mathrm{Eis}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Q}))$, whose image under the comparison map (comparison between the Betti cohomology and the deRham cohomology) is the class of the differential form $N^{k}F^{k+2}_{1/N}dw^{\otimes k}d\tau$. Here $F^{k+2}_{1/N}$ is an Eisenstein series of weight $k+2$ that appears in [12]. Via the base-extension, we can consider $\mathrm{Eis}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Q}_{p}))$. This class does not generally lie in the lattice $\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Z}_{p}))$. But for any integer $c>1$ coprime with $2,3,N$, there exists $\prescript{}{c}{\mathrm{Eis}}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{kj}(\mathbb{Z}_{p}))$ such that
@@
$(u_{a}\circ\iota)_{*}\circ\mathrm{CG}^{[k,k,kj]}:\mathrm{H}^{1}(Y_{\mathbb{Q}}(m,mN),T_{2k-2j}(\mathcal{O}_{E}))\to\mathrm{H}^{2}(Y^{*}_{F}(m,m\mathfrak{N}),T_{kk}(\mathcal{O}_{E})).$
@@
This description is convenient to relate this element with special values of the Asai $L$-function. See [21, Lemma 5.4] for the details.
@@
[21, Proposition 5.5] For any integer $r\geq 0$, we have
@@ 4.4. The $p$-adic Asai $L$-function: $p$-ordinary case
[21, Theorem 7.5] For any integer $c>1$ coprime to $6\mathfrak{N}$, there exists a $p$-adic $L$-function
@@
Here and $(\mathrm{e}_{\mathrm{ord},*})$ is the Hida’s ordinary projector. Note that, roughly, by King’s theory of polylogarithms ([13]), $\prescript{}{c}{L^{\mathrm{As}}_{p}}(\Psi)$ is independent of the twist $j$. Thus, integrating $x^{j}\theta(x)$ against $d\prescript{}{c}{L^{\mathrm{As}}_{p}}(\Psi)$ computes $L^{\mathrm{As}}(\Psi,\overline{\theta},j+1)$. In other words, we have the following equality
@@ 5.1. On mixed level locally symmetric spaces $Y^{*}_{F}(m,m\mathfrak{N})$
See [21, Proof of Proposition 4.4]. If we identify $Y^{*}_{F}(m,m\mathfrak{N})^{(t)}$ with $\Gamma^{*}_{F}(m,m\mathfrak{N})\backslash\mathbb{H}_{3}$ via $\begin{pmatrix}t&0\\ 0&1\end{pmatrix}$, the restriction to this component of the right action of $\begin{pmatrix}1&-a\\ 0&1\end{pmatrix}$ on the adelic symmetric space corresponds to the left action of $\begin{pmatrix}1&ta\\ 0&1\end{pmatrix}$ on $\mathbb{H}_{3}$, i.e, the right action of $\begin{pmatrix}1&-ta\\ 0&1\end{pmatrix}$.
@@ 5.2. Cyclotomic twists calculations
See [17, Remark 8.1.2] for a similar situation in the Hilbert modular form setting.
@@
The proof is similar to the proofs of [14, Theorem 6.2.4] and [17, Theorem 8.1.4] in the étale cohomology setting. For the convenience of readers, we will prove this theorem.
@@
Now as in the proof of [21, Proposition 4.4] and by the calculations involving $\prescript{}{c}{\mathcal{Z}}$, if we write $\mathcal{Z}^{[j]}(t)=(\mathrm{proj}_{t})_{*}\left(\iota_{*}(z)\cup\sum_{i=0}^{j}(-1)^{i}(j-i)!e_{1,r}^{[i]}e_{2,r}^{[j-i]}\otimes e_{1,r}^{[j-i]}e_{2,r}^{[i]}\right)\in\mathrm{H}^{2}(Y^{*}_{F}(p^{r},p^{r}\mathfrak{N})^{(t)},T_{jj}(\mathcal{O}_{E}/p^{r}))$, then
@@
This is proved in [17, Lemma 8.2.1]. See also [14, Lemma 6.3.2]. ∎
@@
This is [21, Proposition 5.6]. We will prove this for the convenience of the readers. We know that for any $t\in(\mathbb{Z}/p^{r})^{\times}$, by definition,
@@
On the other hand, from [21, Proposition 5.5], we have
@@
Now using the following commutative diagram (for the local systems on $Y^{*}_{F}$ or modules, see [14, Proposition 5.1.2] for more details),
@@ 5.3. Patching arguments
This proof is similar to the proof of [22, Proposition 3.3.4]. But here we are proving it in the setting of the Betty cohomology and for the algebraic group over $F$ rather than $\mathbb{Q}$. Write $\mathrm{Res}$ for $\mathrm{Res}_{p^{r}}^{p^{jr}}$. From the equation (5.1), and calculations involving $\prescript{}{c}{\mathcal{Z}}_{p^{r},\mathfrak{N},a}(t)$, we have, modulo $p^{jr}$,
@@
The proof is similar to the proof of [22, Theorem 3.3.5].
@@ 6. Interpolation of twists and construction of the $p$-adic distribution
Recall D-D to be the discriminant of $F$. We have assumed the level $\mathfrak{N}$ to be divisible by some integer $\geq 4$. This is automatic if $p$ is unramified in $F$ and $p\geq 5$. ([21, Page 1692]). Note that we have taken $a\in\mathcal{O}_{F}$ such that $a$ generates $\mathcal{O}_{F}/(p\mathcal{O}_{F}+\mathbb{Z})$. Thus take $a=\dfrac{1}{2}(1+\sqrt{-D})$ if $D\equiv-1\mod 4$, and $a=\dfrac{1}{2}\sqrt{-D}$ if $D\equiv 0\mod 4$. Thus $a-a^{\sigma}=\sqrt{-D}$. Recall $E/\mathbb{Q}_{p}$ is a finite extension large enough such that $F$ embeds into $E$ and all Hecke eigenvalues are also in $E$.
@@ 6.1. Polynomial setup
To construct the distribution, i.e., a power series with unbounded coefficients, we use the technique developed by Perrin-Riou ([24, Subsection 1.2]) and by Büyükboduk–Lei ([6, Section 2]). To use these methods, we need the language of polynomials, or rather polynomials modulo $\omega_{r}(X)=(1+X)^{p^{r}}-1$.
@@ 6.2. Construction of the $p$-adic $L$-function at non-ordinary primes
See [24, Lemme 1.2.1, Lemme 1.2.2] and [6, Lemma 2.2, Lemma 2.3] for the details. ∎
@@
(6.4) $P^{\mathrm{triv}}_{1,j}(T)=\left\langle\phi^{*}_{\Psi},\dfrac{(1=p^{-j}(U_{p})^{-1}_{*})}{(a^{\sigma}-a)^{j}j!{k\choose j}^{2}}\prescript{}{c}{\Xi}^{k,j}_{1,\mathfrak{N},a}\right\rangle.$
@@ 6.3. The interpolation property
The theorem then follows from [21, Theorem 7.5]. ∎
@@ 6.4. $p$-adic distributions without $c$
where $\prescript{}{c}{\mathrm{Eis}}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Z}_{p}))$ appearing in the Definition 4.13. See [14, Theorem 4.4.4] for the details. Following this path, we prove:
@@
We follow [21, Proposition 6.7]. Denote $\left(c^{2}-c^{-2k}\epsilon_{\Psi}(c^{-1})\left(\gamma_{0}^{\log_{u}(c)}\right)^{2}\right)$ by $\upsilon_{c}$. For any positive integers $c,d$ coprime to $6Np$, consider
@@
There is a typo in the equation of $L_{p}^{\mathrm{As}}(\Psi)$ appearing in [21, Proposition 6.7]. There should be $\epsilon_{\Psi}(c^{-1})$ instead of $\epsilon_{\Psi}(c)$.
@@ 7.1. Wach modules and logarithmic matrix
Let $X$ be a variable and write $\mathbb{B}^{+}_{\mathrm{rig},\mathbb{Q}_{p}}$ for the ring of power series $f(X)\in\mathbb{Q}_{p}[[X]]$ such that $f$ converges everywhere inside the unit $p$-adic disk. We equip $\mathbb{B}^{+}_{\mathrm{rig},\mathbb{Q}_{p}}$ with the actions by the Frobeneius $\varphi:X\mapsto(1+X)^{p}-1$ and $\sigma:X\mapsto(1+X)^{\chi(\sigma)}-1$ for $\sigma\in\Gamma$, where $\chi$ is the $p$-adic cyclotomic character such that $\chi(\gamma_{0})=u$. Let $\mathbb{B}^{+}_{\mathrm{rig},E}=E\otimes\mathbb{B}^{+}_{\mathrm{rig},\mathbb{Q}_{p}}$ and inside $\mathbb{B}^{+}_{\mathrm{rig},E}$ there is a subring $\mathbb{A}^{+}_{E}=\mathcal{O}_{E}[[X]]$ which is also equipped with the actions of $\varphi$ and $\Gamma$. Note that there exists a $\Lambda_{E}(\Gamma)$-module isomorphism between $(\mathbb{B}^{+}_{\mathrm{rig},E})^{\psi=0}$ and $\mathcal{H}_{E}(\Gamma)$ called the Mellin transform, where $\psi$ is a left inverse of $\varphi$ such that $\varphi\circ\psi(f(X))=\frac{1}{p}\sum_{\zeta^{p}=1}f(\zeta(1+X)-1)$. Moreover, we can identify $(\mathbb{A}^{+}_{E})^{\psi=0}$ with $\Lambda_{\mathcal{O}_{E}}(\Gamma)$. See [24, B.2.8] for more details.
@@
Note that $\underline{M}$ is unique and in fact $\underline{M}\in M_{2,2}(\mathcal{H}_{E}(\Gamma_{1}))$. See [9, Section 4] and [16] for the details about the construction of $\underline{M}$.
@@
See [9, Proposition 5.2, Lemma 5.2, Lemma5.3, Lemma 5.4] for the details. Note that in [9], $k\geq 2$. Here, we are taking $k\geq 0$, but otherwise, everything is identical. ∎
@@ 7.2. Decomposition of the distribution into measures
For more details about the $p$-stabilizations of Bianchi modular forms, we refer to [23, Section 3.3].
@@
where $c_{\theta,j}$ is independent of $\dagger$, the theorem follows from [9, Theorem 5.5]. ∎
@@
where $\alpha_{\mathfrak{q}},\beta_{\mathfrak{q}}$ are the roots of polynomial $X^{2}-a_{\mathfrak{q}}X+p^{k+1}$, for $\mathfrak{q}\in\{\mathfrak{p},\overline{\mathfrak{p}}\}$. For Galois representation theoretic interpretation, see [10, Section 4].

### v3

@@ 4.2.1. Another description of weight $2$ Asai-Eisenstein elements
[17, Lemma 4.5] The pushforward of $\prescript{}{c}{C}_{m^{2}N}$ along the map $Y_{\mathbb{Q},1}(m^{2}N)\to Y_{\mathbb{Q}}(m,mN),$ given by $z\mapsto mz$ on $\mathbb{H}$, is $\prescript{}{c}{C}_{mN}$.
@@
[17, Proposition 4.4] Let $a\in\mathcal{O}_{F}$ be a generator of $\mathcal{O}_{F}/(p\mathcal{O}_{F}+\mathbb{Z})$. We have
@@ 4.3.2. Relation to the weight $2k$ Eisenstein series
From [17, Proposition 5.2], for an integer $k\geq 0$, we know that there exists a class $\mathrm{Eis}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Q}))$, whose image under the comparison map (comparison between the Betti cohomology and the deRham cohomology) is the class of the differential form $N^{k}F^{k+2}_{1/N}dw^{\otimes k}d\tau$. Here $F^{k+2}_{1/N}$ is an Eisenstein series of weight $k+2$ that appears in [11]. Via the base-extension, we can consider $\mathrm{Eis}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Q}_{p}))$. This class does not generally lie in the lattice $\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Z}_{p}))$. But for any integer $c>1$ coprime with $2,3,N$, there exists $\prescript{}{c}{\mathrm{Eis}}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{kj}(\mathbb{Z}_{p}))$ such that
@@
$(u_{a}\circ\iota)_{*}\circ\mathrm{CG}^{[k,k,j]}:\mathrm{H}^{1}(Y_{\mathbb{Q}}(m,mN),T_{2k-2j}(\mathcal{O}_{E}))\to\mathrm{H}^{2}(Y^{*}_{F}(m,m\mathfrak{N}),T_{kk}(\mathcal{O}_{E})).$
@@
This description is convenient to relate this element with special values of the Asai $L$-function. See [17, Lemma 5.4] for the details.
@@
[17, Proposition 5.5] For any integer $r\geq 0$, we have
@@ 4.4. The $p$-adic Asai $L$-function: $p$-ordinary case
[17, Theorem 7.5] For any integer $c>1$ coprime to $6\mathfrak{N}$, there exists a $p$-adic $L$-function
@@
Here $(\mathrm{e}_{\mathrm{ord},*})$ is the Hida’s ordinary projector. Note that, roughly, by King’s theory of polylogarithms ([13]), $\prescript{}{c}{L^{\mathrm{As}}_{p}}(\Psi)$ is independent of the twist $j$. Thus, integrating $x^{j}\theta(x)$ against $d\prescript{}{c}{L^{\mathrm{As}}_{p}}(\Psi)$ computes $L^{\mathrm{As}}(\Psi,\overline{\theta},j+1)$. In other words, we have the following equality
@@ 5.1. On mixed level locally symmetric spaces $Y^{*}_{F}(m,m\mathfrak{N})$
See [17, Proof of Proposition 4.4]. If we identify $Y^{*}_{F}(m,m\mathfrak{N})^{(t)}$ with $\Gamma^{*}_{F}(m,m\mathfrak{N})\backslash\mathbb{H}_{3}$ via $\begin{pmatrix}t&0\\ 0&1\end{pmatrix}$, the restriction to this component of the right action of $\begin{pmatrix}1&-a\\ 0&1\end{pmatrix}$ on the adelic symmetric space corresponds to the left action of $\begin{pmatrix}1&ta\\ 0&1\end{pmatrix}$ on $\mathbb{H}_{3}$, i.e, the right action of $\begin{pmatrix}1&-ta\\ 0&1\end{pmatrix}$.
@@ 5.2. Cyclotomic twists calculations
See [16, Remark 8.1.2] for a similar situation in the Hilbert modular form setting.
@@
The proof is similar to the proofs of [12, Theorem 6.2.4] and [16, Theorem 8.1.4] in the étale cohomology setting. For the convenience of readers, we will prove this theorem.
@@
Now as in the proof of [17, Proposition 4.4] and by the calculations involving $\prescript{}{c}{\mathcal{Z}}$, if we write $\mathcal{Z}^{[j]}(t)=(\mathrm{proj}_{t})_{*}\left(\iota_{*}(z)\cup\sum_{i=0}^{j}(-1)^{i}(j-i)!e_{1,r}^{[i]}e_{2,r}^{[j-i]}\otimes e_{1,r}^{[j-i]}e_{2,r}^{[i]}\right)\in\mathrm{H}^{2}(Y^{*}_{F}(p^{r},p^{r}\mathfrak{N})^{(t)},T_{jj}(\mathcal{O}_{E}/p^{r}))$, then
@@
This is proved in [16, Lemma 8.2.1]. See also [12, Lemma 6.3.2]. ∎
@@
This is [17, Proposition 5.6]. We will prove this for the convenience of the readers. We know that for any $t\in(\mathbb{Z}/p^{r})^{\times}$, by definition,
@@
On the other hand, from [17, Proposition 5.5], we have
@@
Now using the following commutative diagram (for the local systems on $Y^{*}_{F}$ or modules, see [12, Proposition 5.1.2] for more details),
@@ 5.3. Patching arguments
This proof is similar to the proof of [18, Proposition 3.3.4]. But here we are proving it in the setting of the Betti cohomology and for the algebraic group over $F$ rather than $\mathbb{Q}$. Write $\mathrm{Res}$ for $\mathrm{Res}_{p^{r}}^{p^{jr}}$. From the equation (5.1), and calculations involving $\prescript{}{c}{\mathcal{Z}}_{p^{r},\mathfrak{N},a}(t)$, we have, modulo $p^{jr}$,
@@
The proof is similar to the proof of [18, Theorem 3.3.5].
@@ 6. Interpolation of twists and construction of the $p$-adic distribution
Recall D-D to be the discriminant of $F$. We have assumed the level $\mathfrak{N}$ to be divisible by some integer $\geq 4$. This is automatic if $p$ is unramified in $F$ and $p\geq 5$. ([17, Page 1692]). Note that we have taken $a\in\mathcal{O}_{F}$ such that $a$ generates $\mathcal{O}_{F}/(p\mathcal{O}_{F}+\mathbb{Z})$. Thus take $a=\dfrac{1}{2}(1+\sqrt{-D})$ if $D\equiv-1\mod 4$, and $a=\dfrac{1}{2}\sqrt{-D}$ if $D\equiv 0\mod 4$. Thus $a-a^{\sigma}=\sqrt{-D}$. Recall $E/\mathbb{Q}_{p}$ is a finite extension large enough such that $F$ embeds into $E$ and all Hecke eigenvalues are also in $E$.
@@ 6.1. Polynomial setup
To construct the distribution, i.e., a power series with unbounded coefficients, we use the technique developed by Perrin-Riou ([20, Subsection 1.2]) and by Büyükboduk–Lei ([5, Section 2]). To use these methods, we need the language of polynomials, or rather polynomials modulo $\omega_{r}(X)=(1+X)^{p^{r}}-1$.
@@ 6.2. Construction of the $p$-adic $L$-function at non-ordinary primes
See [20, Lemme 1.2.1, Lemme 1.2.2] and [5, Lemma 2.2, Lemma 2.3] for the details. ∎
@@
(6.4) $P^{\mathrm{triv}}_{1,j}(T)=\left\langle\phi^{*}_{\Psi},\dfrac{(1-p^{-j}(U_{p})^{-1}_{*})}{(a^{\sigma}-a)^{j}j!{k\choose j}^{2}}\prescript{}{c}{\Xi}^{k,j}_{1,\mathfrak{N},a}\right\rangle.$
@@ 6.3. The interpolation property
The theorem then follows from [17, Theorem 7.5]. ∎
@@ 6.4. $p$-adic distributions without $c$
where $\prescript{}{c}{\mathrm{Eis}}^{k}_{N}\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),T_{k}(\mathbb{Z}_{p}))$ appearing in the Definition 4.13. See [12, Theorem 4.4.4] for the details. Following this path, we prove:
@@
We follow [17, Proposition 6.7]. Denote $\left(c^{2}-c^{-2k}\epsilon_{\Psi}(c^{-1})\left(\gamma_{0}^{\log_{u}(c)}\right)^{2}\right)$ by $\upsilon_{c}$. For any positive integers $c,d$ coprime to $6Np$, consider
@@
There is a typo in the equation of $L_{p}^{\mathrm{As}}(\Psi)$ appearing in [17, Proposition 6.7]. There should be $\epsilon_{\Psi}(c^{-1})$ instead of $\epsilon_{\Psi}(c)$.
@@ 7.1. Wach modules and logarithmic matrix
Let $X$ be a variable and write $\mathbb{B}^{+}_{\mathrm{rig},\mathbb{Q}_{p}}$ for the ring of power series $f(X)\in\mathbb{Q}_{p}[[X]]$ such that $f$ converges everywhere inside the unit $p$-adic disk. We equip $\mathbb{B}^{+}_{\mathrm{rig},\mathbb{Q}_{p}}$ with the actions by the Frobeneius $\varphi:X\mapsto(1+X)^{p}-1$ and $\sigma:X\mapsto(1+X)^{\chi(\sigma)}-1$ for $\sigma\in\Gamma$, where $\chi$ is the $p$-adic cyclotomic character such that $\chi(\gamma_{0})=u$. Let $\mathbb{B}^{+}_{\mathrm{rig},E}=E\otimes\mathbb{B}^{+}_{\mathrm{rig},\mathbb{Q}_{p}}$ and inside $\mathbb{B}^{+}_{\mathrm{rig},E}$ there is a subring $\mathbb{A}^{+}_{E}=\mathcal{O}_{E}[[X]]$ which is also equipped with the actions of $\varphi$ and $\Gamma$. Note that there exists a $\Lambda_{E}(\Gamma)$-module isomorphism between $(\mathbb{B}^{+}_{\mathrm{rig},E})^{\psi=0}$ and $\mathcal{H}_{E}(\Gamma)$ called the Mellin transform, where $\psi$ is a left inverse of $\varphi$ such that $\varphi\circ\psi(f(X))=\frac{1}{p}\sum_{\zeta^{p}=1}f(\zeta(1+X)-1)$. Moreover, we can identify $(\mathbb{A}^{+}_{E})^{\psi=0}$ with $\Lambda_{\mathcal{O}_{E}}(\Gamma)$. See [20, B.2.8] for more details.
@@
Note that $\underline{M}$ is unique and in fact $\underline{M}\in M_{2,2}(\mathcal{H}_{E}(\Gamma_{1}))$. See [8, Section 4] and [14] for the details about the construction of $\underline{M}$.
@@
See [8, Proposition 5.2, Lemma 5.2, Lemma5.3, Lemma 5.4] for the details. Note that in [8], $k\geq 2$. Here, we are taking $k\geq 0$, but otherwise, everything is identical. ∎
@@ 7.2. Decomposition of the distribution into measures
For more details about the $p$-stabilizations of Bianchi modular forms, we refer to [19, Section 3.3].
@@
where $c_{\theta,j}$ is independent of $\dagger$, the theorem follows from [8, Theorem 5.5]. ∎
@@
where $\alpha_{\mathfrak{q}},\beta_{\mathfrak{q}}$ are the roots of polynomial $X^{2}-a_{\mathfrak{q}}X+p^{k+1}$, for $\mathfrak{q}\in\{\mathfrak{p},\overline{\mathfrak{p}}\}$. For Galois representation theoretic interpretation, see [9, Section 4].